[Hallc_running] [Revised Logentry] RC Daily Update - RC Meeting today @4:15pm

bduran at jlab.org bduran at jlab.org
Sat Aug 29 15:40:01 EDT 2020

Logentry Text:
Yesterday, the accelerator was back up around 9pm. 

Earlier around 5:30pm, an NMR measurement (showed only 13% polarization) and noisy laser spectrum made us to find out that accidentally the longitudinal lasers were turned on instead of transverse. At that moment, we turned the longitudinal lasers off and turned transverse on and started pumping in transverse direction. 
When beam came back we proceeded with the reference cell (He3) running at Kin B and Kin Y (shms_p=-6.4, hms_p=-4.2, shms_theta=14.5, hms_theta=16.4) as the target polarization was 26.7%. After 4 hours of reference cell (He3) running, we evacuated the cell and ran about 4 hours with reference cell vacuum at the same kinematics. We kept running at this configuration until 8:25am when another NMR measurement is performed --> 36%. We switched back to Kin C and Kin Z (shms_p=-5.6, hms_p=-4., shms_theta=18.0, hms_theta=20.0) for polhe3 transverse data taking. The current plan is to stay here until at least mid swing shift. We are exploring the longitudinal data taking option later during late swing shift to attempt to understand the relatively low transverse polarization better.

HMS Q3 threw an interlock at 11:20am (https://logbooks.jlab.org/entry/3837011). Although reset was successful, it wouldn't ramp up. The magnet GUI was showing "lock" in red instead of "remote". After reaching out Amy, we took a hall access at 1:51pm when Jack cycled the power supply. Xiaochao ramped the Q3 up for -4.00 GeV. The magnet is in its happy state now (https://logbooks.jlab.org/entry/3837084).

NMR Measurements:

@9:45pm: 26.7% https://logbooks.jlab.org/entry/3836613

@12:00am: 30.1% https://logbooks.jlab.org/entry/3836686

@2:00am: 32.5% https://logbooks.jlab.org/entry/3836727

@06:14am: 35.5%  https://logbooks.jlab.org/entry/3836822

@08:25: 36.% https://logbooks.jlab.org/entry/3836884

@1:57pm: 36.8% https://logbooks.jlab.org/entry/3837069

Accelerator status: After beam was back, MCC struggled to keep the beam stable during the swing and owl shift due to bleedthrough of C beam to A: https://logbooks.jlab.org/entry/3836627,  https://logbooks.jlab.org/entry/3836633, https://logbooks.jlab.org/entry/3836636 . At today's MCC meeting, the steps need to be taken to troubleshoot this issue were discussed (https://logbooks.jlab.org/entry/3836933, https://logbooks.jlab.org/entry/3836985, https://logbooks.jlab.org/entry/3836644). The issue is pretty much resolved around 11:30am (https://logbooks.jlab.org/entry/3836994). 

Here is the beam status in Hall C for the past 4 days: https://logbooks.jlab.org/entry/3837008

Starting 8am on Monday morning there will be a pass change (5 pass --> 1 pass). We will have about 4 hours of down time. We will talk about the tasks will be performed during this down time and 1 pass running plans at today's RC Meeting. See the connection info below:

Bluejeans remote connection information:
To join the Meeting:
To join via Room System:
Video Conferencing System: bjn.vc -or-
Meeting ID : 861439920
To join via phone :
1) Dial: +1.408.740.7256 (US (San Jose))
+1.888.240.2560 (US Toll Free))
+1.408.317.9253 (US (Primary, San Jose))
(see all numbers - http://bluejeans.com/numbers)
2) Enter Conference ID : 861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3837106
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200829/d52df8a0/attachment-0001.html>

More information about the Hallc_running mailing list